lazytaz:

Nicole invests a total of $5000 into two different investments. One investment is a GIC that earns 5% interest. The other investment is a savings plan that earns 3% interest. At the end of the year, Nicole earned $220 in interest. Determine how much money Nicole put in each investment.

jhonyy9:

you need to solve this sistem for x and y x +y = 5000 0,05x +0,03y = 220 @Vocaloid is this true ? pls.

jhonyy9:

ty

lazytaz:

I don't quite understand? can you explain how to get the x and y?

jhonyy9:

you can use substitution or elimination method to solve this sistem @Laylalyssa

lazytaz:

but what exactly am i substituting?

Laylalyssa:

yes. both would get the same answer

Laylalyssa:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @lazytaz but what exactly am i substituting? \(\color{#0cbb34}{\text{End of Quote}}\) the variable I believe

lazytaz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Laylalyssa \(\color{#0cbb34}{\text{Originally Posted by}}\) @lazytaz but what exactly am i substituting? \(\color{#0cbb34}{\text{End of Quote}}\) the variable I believe \(\color{#0cbb34}{\text{End of Quote}}\) which one is the variable $5000, 5%, 3%, or $220?

Laylalyssa:

@jhonyy9

jhonyy9:

no x +y = 5000 => x = 5000 -y and this value of x substitute in place of x inside second equation and calcule the y value 0,05x +0,03y = 220